# Two prover perfect zero knowledge for MIP*

### v1

@@ 8. Perfect zero knowledge
Suppose $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ is a BCS, and let $\operatorname{Tab}(B)=(Y,\{(W_{i},D_{i})\}_{i=1}^{m})$. Then there is a classical homomorphism $\alpha:\mathcal{A}(B)\to\mathcal{A}(\operatorname{Tab}(B))$ such that $\alpha(\sigma_{i}(x))=\sigma_{i}(x)$ for all $i\in[m]$ and $x\in V_{i}$.
Conversely, let $R=\{r_{i}(j,k):(i,j,k)\in[m]\times[2]\times[d_{i}-1]\}$ be the set of randomizers in $Y$. If $r\in S_{5}^{R}$, then there is a classical homomorphism $\beta_{r}:\mathcal{A}(\operatorname{Tab}(B))\to\mathcal{A}(B)$ such that $\beta_{r}(\sigma_{i}(x))=\sigma_{i}(x)$ for all $i\in[m]$ and $x\in V_{i}$, and $\beta_{r}(x)=e^{2\pi ij/120}$ for all $x\in R$, where $r(x)=\gamma_{j}$ in the enumeration of $S_{5}$ fixed above.

### v2

@@ 9. Perfect zero knowledge
Let $f:S_{5}^{m}\to S_{5}$ be a function, and suppose $(V,C)$ is a boolean constraint encoding the constraint $x=f(y_{1},\ldots,y_{m})$ on permutation-valued variables $x,y_{1},\ldots,y_{m}$. If $1201\leq n<$ 120, then
$x^{n}=\sum_{a}c_{a}y_{1}^{a_{1}}\cdots y_{m}^{a_{m}}$
for some coefficients $c_{a}\in\mathbb{C}$, where the sum is over all integer vectors $a=(a_{1},\ldots,a_{m})$ with $1200\leq a_{1},\ldots,a_{m}<$ 120. Furthermore, if for every $\pi_{1},\ldots,\pi_{m-1}\in S_{5}$, the set $\{f(\pi_{1},\ldots,\pi_{k-1},\pi,\pi_{k},\ldots,\pi_{m-1}):\pi\in S_{5}\}$ is equal to $S_{5}$, then $c_{a}=0$ if $a_{k}=0$.